OpenStudy (anonymous):

Yuri drove his car to the train station and then traveled six times as far by train to get to work. He traveled a total of 42 miles to work. How far did he drive?

OpenStudy (jamesj):

Let x be the distance travelled by car. Then what is the distance he travelled by train? Call it y. What is y as a function of x? Then solve x + y = 42 for x

OpenStudy (anonymous):

these are the answers available??? 5 miles 6 miles 7 miles 8 miles

OpenStudy (jamesj):

y = distance travelled by train = 6 times distance travelled by car. If distance travelled by = x, what is y = ...?

OpenStudy (anonymous):

I dont get this one

OpenStudy (jamesj):

(distance travelled by car) + (distance travelled by train) = 42 x + y = 42 Now we want an equation in one variable, x, the distance travelled by car. What is distance travelled by train as a function of the distance travelled by car? What does the problem say?

OpenStudy (jamesj):

The problem says he "traveled six times as far by train". Hence what is y as a function of x? y = ...what... times x?

OpenStudy (anonymous):

idk!!

OpenStudy (hang254):

just look at it as another algerbraic equation 6x=42

OpenStudy (jamesj):

He travels six times as far by train. Hence y = 6x. Therefore if x + y = 42 x + 6x = 42 Now solve for x.

OpenStudy (jamesj):

@hang254, no, not quite.

OpenStudy (hang254):

but, you would get the same answer....

OpenStudy (jamesj):

No 6x = 42, isn't right. x + 6x = 42 hence 7x = 42

OpenStudy (hang254):

oh well